Black hair color in rabbits is due to the inherita
Answer #1

Ans. 31. Genotype of the parents = Bb

Gametes produced by each Parent = B, b

Probability of fusion of a particular gamete = No. of the specific gamete/ total no. of gametes

Or, p(b)= (1/2)

Genotype of white bunny = bb; i.e. from fusion of two ‘b’ gametes.

Each ‘b’ gamete has a (1/2) probability of fusion with another gamete irrespective of the number of progeny [i.e. 3 is less than 4- the maximum theoretical value of the progeny produced in a (B / b) X (b/ b)]. Production of 3 progenies does not eliminate the probability of the 4th progeny, which has an equal probability of appearing in the 3 litters group.

Thus, the maximum probability of a bb (white) bunny =

                                                                        Probability of b (male) x probability of b (female)

                                                                        = (1/2) x (1/2) = 1/4 = 0.25

Correct option – D. 0.25

32. a. Probability of the ‘first being white’ bb = p (b) x p (b) = 1/4

b. Probability of the second bunny being black.

Possible genotypes of black bunnies = BB, Bb

Calculate probability using (B/ b) x (B/ b) as follow-

            P (BB) = p (B, male) x p (B, female) = (1/2) x (1/2) = (1/4)

            P (Bb) = p (B, male) x p (b, female) + p (b, male) x p (B, female) = (1/4) + (1/4) = 1/2

Total probability of getting a black bunny = p (BB) + p (Bb)

                                                                    = (1/4) + (1/2) = (3/4)

c. Probability of getting the third bunny as white.

Theoretically only 1 white bunny can be produced per generation. That the parent bunnies must to another round of reproduction to produce another white bunny. In summary, the parents shall go 2nd rounds of reproduction to get the second white bunny. Each generation the probability of getting a white bunny = (1/4)

i.e. probability of getting a white bunny in next generation is also (1/4)

Now, the maximum probability of getting 1st white, 2nd black and 3rd white bunny in a group of 3 litters = p (a) x p (b) x p (c) = (1/4) x (3/4) x (1/4) = 3/ 64

Correct option – A. 3/64